OpenStudy (jiteshmeghwal9):

Plz give me a tutrial on perfect square method for quadratic equation :)

hartnn (hartnn):

Lets take an example : x^2 + 6x+8 =0

hartnn (hartnn):

Step 1 : what is the co-efficient of 'x' ?

OpenStudy (jiteshmeghwal9):

6

hartnn (hartnn):

step 1 : Divide it by 2.

hartnn (hartnn):

step 2 *

OpenStudy (jiteshmeghwal9):

3x

hartnn (hartnn):

step 3 :Square it.

hartnn (hartnn):

not 3x, just 3

OpenStudy (jiteshmeghwal9):

9x

hartnn (hartnn):

where are 'x's coming from ???

hartnn (hartnn):

the co-efficient of 'x' is just 6

OpenStudy (jiteshmeghwal9):

ok sorry

OpenStudy (jiteshmeghwal9):

it's just 9

hartnn (hartnn):

step 4 : Add and subtract 9 .

OpenStudy (jiteshmeghwal9):

To whom i add & subtract 9

OpenStudy (jiteshmeghwal9):

I mean which number

hartnn (hartnn):

\(x^2+6x+9 -9+8=0 \\ x^2+6x+9=9-8\) get this....

OpenStudy (jiteshmeghwal9):

yup

hartnn (hartnn):

now lefty is a perfect square.

OpenStudy (jiteshmeghwal9):

mean left side equation

hartnn (hartnn):

lol, yes!

OpenStudy (jiteshmeghwal9):

Ok! thanx i gt it :)

hartnn (hartnn):

want to solve for general quadratic ? \(ax^2+bx+c=0\)

OpenStudy (jiteshmeghwal9):

no it's example is given my book

OpenStudy (jiteshmeghwal9):

in*

hartnn (hartnn):

okk...

OpenStudy (unklerhaukus):

Another example : \[x^2 + 4x+5 =0 \]\[\rightarrow (x^2 + 4x+2)+3 =0 \]\[\rightarrow (x +2)^2+3 =0 \]

hartnn (hartnn):

you mean (x+2)^2+1 ??

OpenStudy (unklerhaukus):

ah yes, how did you mess that up.

OpenStudy (unklerhaukus):

\[{\color{red}{**}}\]\[x^2+4x+5=0\]\[→(x^2+4x+4)+1=0\]\[→(x+2)^2+1=0\]